# Krait: A Backdoor Attack Against Graph Prompt Tuning

### v1

@@ III-A Metric Design
In this subsection, we introduce two node-centric homophily metrics and three centroid similarity metrics to investigate the behaviors of graph prompt tuning before and after backdoor injections.
@@ III-A1 Node-centric Homophily
Generally, homophily describes the phenomenon where connected nodes tend to share similar labels or features [unnoticeability_node_centric_homo]. Since many real graphs are homophilic and most existing GNNs implicitly follow homophily assumptions, we leverage homophily to investigate the behaviors of graph prompt tuning before and after backdoor injections.
However, graph prompt tuning generally constructs an ego-network for each targeted node, which involves transforming nodes into local subgraphs. Each ego-network is paired with a specific node label, but the nodes within the ego-network do not have labels, which is also the case with the graph prompt appended to each ego-network. Therefore, we cannot directly calculate label-based homophily. To address this issue, based on node-centric homophily [unnoticeability_node_centric_homo], we propose two metrics—local-subgraph homophily and global-view homophily—to compute node feature similarity between a node and its $1$-hop neighbors in different ways.

### v2

@@ III-A Metric Design
We first introduce two node-centric homophily metrics and three centroid similarity metrics to investigate the behaviors of graph prompt tuning before and after backdoor injections.
@@ III-A1 Node-centric Homophily
Homophily describes the phenomenon where connected nodes tend to share similar labels or features [2]. Since many real graphs are homophilic and most existing GNNs implicitly follow homophily assumptions, we leverage homophily to investigate the behaviors of graph prompt tuning before and after backdoor injections.
However, graph prompt tuning generally constructs an ego-network for each targeted node, which involves transforming nodes into local subgraphs. Each ego-network is paired with a specific node label, but the nodes within the ego-network do not have labels, which is also the case with the graph prompt appended to each ego-network. Therefore, we cannot directly calculate label-based homophily. To address this issue, based on node-centric homophily [2], we propose two metrics—local-subgraph homophily and global-view homophily—to compute node feature similarity between a node and its $1$-hop neighbors in different ways.